Calculate the minimum thickness of a soap bubble film (n=1.33) that will result in constructive interference in the reflected light if the film is illuminated by light of wavelength 602nm

Grade:12

1 Answers

Nirmal Singh.
10 years ago
for constructive interference
2t = (m + 1/2 ) Lambda / n
t = (m + 1/2 ) Lambda /2 n
t = (0+1/2) * 602 / 2*1.33 where m = 0
t = 113.16 nm
Thanks & Regards,
Nirmal Singh
Askiitians Faculty